# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ Appendix C Linear Program for TBSM when PTC is satisfied exactly
This function is indeed, an indicator function of whether the outputs $\bm{x}$ and tokens $\bm{t}$ are consistent. Note that clearly for any $\bm{t}$ there is only one consistent $\bm{x}$, and more interestingly for any PTC tuple $\bm{x}=(a,b,c)$, i.e. $a\oplus b\oplus c=1$, there are exactly two consistent tokens $\bm{t}$, which are flipped of each other. For example for $\bm{x}=(1,1,1)$ we find $\bm{t}=(1,1,1)$, or $\bm{t}=(0,0,0)$.
@@
Note that as mentioned before, for any PTC tuple of $\bm{x}_{1}$, there are exactly two consistent $\bm{t}$, therefore the sum in $(i)$, is only over two terms.
Dephasing noise: the local dephasing noise given by the channel $\mathcal{E}_{d}:\outerproduct{\psi}{\psi}\mapsto\varrho$ with
@@
where |$\ket{\psi}=\lambda_{0}\ket{01}+\lambda_{1}\ket{10}$. Note that the resulting distribution $P$ in the presence of the dephasing noise is still a PTC distribution and can be written as
@@
Where $P_{Q}$ is the quantum distribution obtained without the presence of dephasing noise($d=0$), and $P_{C}$ is the classical distribution($d=1$), therefore we have, $P=P_{Q}+(P_{C}-P_{Q})e$ for $e:=(1-(1-d)^{3})$ and using (107) finding the minimum amount of dephasing noise $d$ for which the distribution becomes local can be writen as
Minimize $\text{Minimize \ }e$ (109)
@@
Solving this LP, we find the optimum value $e^{*}$, which translates to the optimum dephasing noise $d^{*}=1+\sqrt[3]{e^{*}-1}$. We used MOSEK Optimizer API for Python to solve the Linear Program [32].
@@ Appendix D Linear Program for noisy TBSM distributions
For probability distributions $p(x)$ and $q(x)$ the following statements are equivalent,
$(i)\ \text{There exist probability distributions $p^{\prime}$ and$q^{\prime}$ such that }p(x)=q(x)+\varepsilon p^{\prime}(x)-\varepsilon q^{\prime}(x)$
@@
$(ii)\Rightarrow(i):$ let’s define $r(x)=p(x)-q(x)$, then we have $\sum_{x}r(x)=0$ and $\sum_{x}|r(x)|\leq\varepsilon$ we define $p^{\prime}(x)=\frac{|r(x)|}{\sum_{x}|r(x)|}$ and $q^{\prime}(x)=p^{\prime}(x)-\frac{r(x)}{\varepsilon}$, which are valid probability distributions satisfying (i).
@@
(*$(*^{\prime})\,P(\bm{x}_{1},\bm{x}_{2})=\underbrace{\sum_{\bm{t}}P_{\scriptscriptstyle PTC}(\bm{x}_{1}|\bm{t})P(\bm{x}_{2},\bm{t})}_{\widetilde{P}(\bm{x}_{1}\bm{x}_{2})}-\text{f}_{\varepsilon}P^{\prime}(\bm{x}_{1},\bm{x}_{2})+\text{f}_{\varepsilon}\widetilde{P}^{\prime}(\bm{x}_{1},\bm{x}_{2})$
@@
Note that $P(t_{\alpha},t_{\beta},t_{\gamma})=p_{\alpha}(t_{\alpha})p_{\beta}(t_{\beta})p_{\gamma}(t_{\gamma})$ is implied by the last constraint. Furthermore, this constraint is coming from the network structure which implies $P(a_{2}|0,t_{\beta},t_{\gamma})=P(a_{2}|1,t_{\beta},t_{\gamma})$. Moreover, note that here $p_{\alpha}(t),p_{\beta}(t),p_{\gamma}(t)$ correspond to the PTC distribution $\widetilde{P}$ and not the observed distribution $P$, therefore we do not know their values and we consider them as additional variables. Note that the quadratic constrains in $(iii)$ are not convex, i.e. if we write these constrains in the form of $x^{T}Qx$, the matrix $Q$ is not positive semidefinite.
@@
First notice that $p_{\alpha}(t_{\alpha}=0)=\frac{1}{2}\big{(}1+\sqrt{|\frac{\widetilde{E}_{b_{1}}\widetilde{E}_{c_{1}}}{\widetilde{E}_{a_{1}}}|}\big{)}$. Where, $\widetilde{E}_{x_{1}}=\widetilde{P}(x_{1}=0)-\widetilde{P}(x_{1}=1)$. It’s straightforward to see that the total variation distance leads to the conclusion that $|\widetilde{E}_{x_{1}}-E_{x_{1}}|\leq 2\text{f}_{\varepsilon}$. In fact, a slightly better bound can be obtained by using the fact that $\widetilde{P}$ is in the PTC slice, see section D.2 below,
@@ D.1 Collection of LPs (the grid technique for $p_{\xi}(t_{\xi})$)
Finally, note that the LP has to be satisfied for some values $p_{\xi}(t_{\xi})$ of the token probabilities in the intervals $[l_{\alpha},u_{\alpha}]$. In practice, it is convenient to divide each interval into smaller intervals $[l_{\alpha}+k\frac{u_{\alpha}-l_{\alpha}}{M},l_{\alpha}+(k+1)\frac{u_{\alpha}-l_{\alpha}}{M}]$, and verify that no solution can be found within any of the $M^{3}$ possible combinations of intervals. This gives a better result as compared to the harsher relaxation in Eqs. (116,117). In particular, for the figures 5,6,10,11, and 12 we used $M=16$, while to obtain the best bounds $\omega=0.544\%,\omega^{\prime}=0.504\%$ white noise and $\varepsilon=0.24\%$ total-variation distance we went to $M=128$.
@@ D.2 Development of a tighter bound on $\widetilde{E}_{x_{1}}$
$P(\bm{x}_{1},\bm{\xi}\in\Lambda)=P(\bm{x}_{1},\bm{\xi}\in\Upsilon)+(\text{f}_{\varepsilon}-\varepsilon)P^{\prime\prime}(\bm{x}_{1})$
$\widetilde{P}(\bm{x}_{1})=P(\bm{x}_{1},\bm{\xi}\in\Upsilon)+\text{f}_{\varepsilon}\widetilde{P^{\prime}}(\bm{x}_{1})$
@@ Appendix E Other noise models
$E_{x_{1},x_{2}}=(1-\omega^{\prime})\outerproduct{\phi_{x_{1},x_{2}}}{\phi_{x_{1},x_{2}}}+\omega^{\prime}\frac{\mathds{1}}{2}\otimes\frac{\mathds{1}}{2}.$ (118)
Note that here for the resulting distribution we have $\text{Pr}(a_{1}\oplus b_{1}\oplus c_{1}=1)=\frac{1}{2}\left(1+(1-\omega^{\prime})^{3}\right)$. In Fig.  11(a), we depict the robustness to measurement white noise of the TBSM distributions with $\lambda_{0}^{2}=0.22$. The maximal amount of noise for which we can prove that nonlocality of the distribution is given by $\omega^{\prime}=0.504\%$ for the case $\phi_{u}=0.42$ and $\phi_{w}=0$.
@@
(a) White noise at the masurement
@@
$E_{x_{1},x_{2}}$ $=(1-p)\outerproduct{\phi_{x_{1},x_{2}}}{\phi_{x_{1},x_{2}}}\qquad\forall(x_{1},x_{2})\neq(1,0).$ (122)
For this strategy, we prove the nonlocality of the resulting distribution for up to a no-click probability of $p=0.7\%$, see Fig.  12(a).
Note that if e.g. the sources produce two photons entangled in polarization |$\ket{0}=\ket{1_{h}}$ and |$\ket{1}=\ket{1_{v}}$, a limited photon transmission (or detector efficiency) $\eta$ would correspond to the no-click noise with $(1-p)=\eta^{2}$.
@@
For $\phi_{w}=0$ the TBSM distributions (RGB4) can be realised with single-photon entanglement (vacuum state |$\ket{0}$ and single photon |$\ket{1}$), linear optics and photon number resolving detectors [28]. In this case the dominant noise is the limited efficiency of the detectors, which can be modeled as a photon loss channel with transmission $\eta$. Formally, we have the channel
@@
$K_{0}$ =|$=\outerproduct{0}{0}+\sqrt{\eta}\outerproduct{1}{1}$ | (124)
$K_{1}$ $=\sqrt{1-\eta}\outerproduct{0}{1}$ | (125)
@@
Recall that for $\phi_{w}=0$ the original projectors for $x_{1}=1$ read $\Pi_{00}=\outerproduct{00}{00}$ | and $\Pi_{01}=\outerproduct{11}{11}$ |, while the lossy POVM elements read
@@ Appendix F Linear Program for the total variation distance
We will prove that all distributions in $\mathcal{B}_{\varepsilon}(\bar{P})$ are nonlocal by contradiction. Let us assume that the ball contains some triangle local distribution $P$, then from Result 3 there must exist a local PTC distribution $\widetilde{P}$ with $\delta(\widetilde{P}(\bm{x}_{1},\bm{x}_{2}),P(\bm{x}_{1},\bm{x}_{2}))\leq\text{f}_{\varepsilon}$ and $\delta(\widetilde{P}(\bm{x}_{2}),P(\bm{x}_{2}))=0$. By triangle inequality, $\widetilde{P}$ must also satisfy
@@
$\bar{P}(\bm{x}_{1},\bm{x}_{2})=\underbrace{\sum_{\bm{t}}P_{\scriptscriptstyle PTC}(\bm{x}_{1}|\bm{t})P(\bm{x}_{2},\bm{t})}_{\widetilde{P}(\bm{x}_{1}\bm{x}_{2})}-(\varepsilon+\text{f}_{\varepsilon})\bar{P}^{\prime}(\bm{x}_{1},\bm{x}_{2})+(\varepsilon+\text{f}_{\varepsilon})\widetilde{P}^{\prime}(\bm{x}_{1},\bm{x}_{2})$ (149)
@@
$\bar{P}(\bm{x}_{2})=\widetilde{P}(\bm{x}_{2})-(\varepsilon+\text{f}_{\varepsilon})\bar{P}^{\prime}(\bm{x}_{2})+(\varepsilon+\text{f}_{\varepsilon})\widetilde{P}^{\prime}(\bm{x}_{2})$ (150)
@@
$\delta(\widetilde{P}^{\prime}(\bm{x}_{2}),\bar{P}^{\prime}(\bm{x}_{2}))=\frac{1}{\varepsilon+\text{f}_{\varepsilon}}\delta(\widetilde{P}(\bm{x}_{2}),\bar{P}(\bm{x}_{2}))\leq\frac{\varepsilon}{\varepsilon+\text{f}_{\varepsilon}}:=\alpha$ (151)
Note that as $2.5\varepsilon\leq\text{f}_{\varepsilon}\leq 3\varepsilon$, we have $71/4\leq\alpha\leq 2/$ 7, moreover for small values of $\varepsilon$, we have $\text{f}_{\varepsilon}\approx 2.5\varepsilon$ and $\alpha\approx 2/7$. Again by Lemma 1, we have
$(4.iii^{\prime})\,\sum_{\bm{x}_{1}}\widetilde{P}^{\prime}(\bm{x}_{1},\bm{x}_{2})-\sum_{\bm{x}_{1}}\bar{P}^{\prime}(\bm{x}_{1},\bm{x}_{2})+\alpha\widetilde{P}^{\prime\prime}(\bm{x}_{2})-\alpha\bar{P}^{\prime\prime}(\bm{x}_{2})=0$ (152)
@@
Notice that $p_{\alpha}(t_{\alpha}=0)=\frac{1}{2}\big{(}1+\sqrt{|\frac{\widetilde{E}_{b_{1}}\widetilde{E}_{c_{1}}}{\widetilde{E}_{a_{1}}}|}\big{)}$. Where, $\widetilde{E}_{x_{1}}=\widetilde{P}(x_{1}=0)-\widetilde{P}(x_{1}=1)$. It’s straightforward to see that the total variation distance $\delta(\widetilde{P},\bar{P})\leq\text{f}_{\varepsilon}+\varepsilon$, leads to the conclusion that $|\widetilde{E}_{x_{1}}-\bar{E}_{x_{1}}|\leq 2(\text{f}_{\varepsilon}+\varepsilon)$. Where $\bar{E}_{x_{1}}=\bar{P}(x_{1}=0)-\bar{P}(x_{1}=1)$ is computed from the target PTC distribution $\bar{P}$. Hence, if all the correlators for the target distribution satisfy $|\bar{E}_{x_{1}}|\geq 2(\text{f}_{\varepsilon}+\varepsilon)$ one can use Eq. (5) to obtain
$\begin{split}l_{\alpha}:=\frac{1}{2}\left(1+\sqrt{\frac{(|\bar{E}_{b_{1}}|-2(\text{f}_{\varepsilon}+\varepsilon))(|\bar{E}_{c_{1}}|-2(\text{f}_{\varepsilon}+\varepsilon))}{|\bar{E}_{a_{1}}|+2(\text{f}_{\varepsilon}+\varepsilon)}}\right)\leq\ p_{\alpha}(0)\ \leq\frac{1}{2}\left(1+\sqrt{\frac{(|\bar{E}_{b_{1}}|+2(\text{f}_{\varepsilon}+\varepsilon))(|\bar{E}_{c_{1}}|+2(\text{f}_{\varepsilon}+\varepsilon))}{|\bar{E}_{a_{1}}|-2(\text{f}_{\varepsilon}+\varepsilon)}}\right):=u_{\alpha}\quad\circlearrowleft.\end{split}$ (154)
@@
$\begin{split}\ l_{\alpha}\sum_{b_{2},c_{2}}{P(a_{2},b_{2},c_{2},0,t_{\beta},t_{\gamma})}-(1-l_{\alpha})\sum_{b_{2},c_{2}}P(a_{2},b_{2},c_{2},1,t_{\beta},t_{\gamma})\leq 0\\ \ u_{\alpha}\sum_{b_{2},c_{2}}{P(a_{2},b_{2},c_{2},0,t_{\beta},t_{\gamma})}-(1-u_{\alpha})\sum_{b_{2},c_{2}}P(a_{2},b_{2},c_{2},1,t_{\beta},t_{\gamma})\geq 0\end{split}$ (155)
Combining Equations 149, 152 155, we get a set of linear constraints which should be satisfied if there exists any local distribution inside the ball $\mathcal{B}_{\varepsilon}(\bar{P})$. Therefore, the infeasibility of this LP, proves the non-locality of all of the points in $\mathcal{B}_{\varepsilon}(\bar{P})$. (See Fig.  6). Again to find tighter bounds we use the grid technique.

### v3

@@ Appendix C Linear Program for TBSM when PTC is satisfied exactly
This function is, indeed, an indicator function of whether the outputs $\bm{x}$ and tokens $\bm{t}$ are consistent. Note that clearly for any $\bm{t}$ there is only one consistent $\bm{x}$, and more interestingly for any PTC tuple $\bm{x}=(a,b,c)$, i.e. $a\oplus b\oplus c=1$, there are exactly two consistent tokens $\bm{t}$, which are flipped of each other. For example for $\bm{x}=(1,1,1)$ we find $\bm{t}=(1,1,1)$, or $\bm{t}=(0,0,0)$.
@@
Note that as mentioned before, for any PTC tuple of $\bm{x}_{1}$, there are exactly two consistent $\bm{t}$; therefore the sum in $(i)$, is only over two terms.
@@
[…]
@@
Where $P_{Q}$ is the quantum distribution obtained without the presence of dephasing noise($d=0$), and $P_{C}$ is the classical distribution($d=1$), therefore we have, $P=P_{Q}+(P_{C}-P_{Q})e$ for $e:=(1-(1-d)^{3})$ and using (103) finding the minimum amount of dephasing noise $d$ for which the distribution becomes local can be written as
$\text{Minimize \ }e$ (105)
@@
Solving this LP, we find the optimum value $e^{*}$, which translates to the optimum dephasing noise $d^{*}=1+\sqrt[3]{e^{*}-1}$. We used the MOSEK Optimizer API for Python to solve the Linear Program [40].
@@ Appendix D Linear Program for noisy TBSM distributions
Let $p(x)$ and $q(x)$ be probability distributions defined on the discrete probability space $(\Omega,\mathcal{F})$. The following statements are equivalent:
[…]
@@
$(ii)\Rightarrow(i):$ let’s define $A\subset\Omega$ as the subset within which $p(x)\geq q(x)$, then $(ii)$ implies that:
$\sum_{x\in A}\big{(}(p(x)-q(x))\big{)}+\sum_{x\in\bar{A}}\big{(}(q(x)-p(x))\big{)}\leq 2\varepsilon$
let’s say $\sum_{x\in A}\big{(}(p(x)-q(x))\big{)}+\sum_{x\in\bar{A}}\big{(}(q(x)-p(x))\big{)}=2\varepsilon^{\prime}$ for $\varepsilon^{\prime}\leq\varepsilon$. Moreover, we have $\sum_{x\in A}\big{(}(p(x)-q(x))\big{)}-\sum_{x\in\bar{A}}\big{(}(q(x)-p(x))\big{)}=\sum_{x\in\Omega}\big{(}(p(x)-q(x))\big{)}=0$, therefore
$\sum_{x\in A}\big{(}(p(x)-q(x))\big{)}=\sum_{x\in\bar{A}}\big{(}(q(x)-p(x))\big{)}=\varepsilon^{\prime}$ (108)
Let’s define $p^{\prime}(x)$ and $q^{\prime}(x)$ as follows:
$p^{\prime}(x)=\begin{cases}\frac{p(x)-q(x)}{\varepsilon}+\frac{\varepsilon-\varepsilon^{\prime}}{\varepsilon}f(x)&x\in A\\ \frac{\varepsilon-\varepsilon^{\prime}}{\varepsilon}f(x)&x\in\bar{A}\end{cases}$
$q^{\prime}(x)=\begin{cases}\frac{\varepsilon-\varepsilon^{\prime}}{\varepsilon}f(x)&x\in A\\ \frac{q(x)-p(x)}{\varepsilon}+\frac{\varepsilon-\varepsilon^{\prime}}{\varepsilon}f(x)&x\in\bar{A}\end{cases}$
for any valid probability distribution $f(x)$ on the same space. It is easy to see that $p^{\prime}(x)$ and $q^{\prime}(x)$ are valid probability distributions satisfying (i). It is also interesting to note that, in the case of equality $\delta_{TV}(p,q)=\varepsilon$, $p^{\prime}(x)$ and $q^{\prime}(x)$ respectively have a non-zero support only on $A$ and $\bar{A}$
Note that equality $\delta_{TV}(p,q)=\varepsilon^{\prime}$ is achieved iff $\varepsilon^{\prime}$ is the maximum $\varepsilon$ satisfying $(i)$ .
@@
$(*^{\prime})\,P(\bm{x}_{1},\bm{x}_{2})=\underbrace{\sum_{\bm{t}}P_{\scriptscriptstyle PTC}(\bm{x}_{1}|\bm{t})P(\bm{x}_{2},\bm{t})}_{\widetilde{P}(\bm{x}_{1}\bm{x}_{2})}-\text{f}_{\varepsilon}P^{\prime}(\bm{x}_{1},\bm{x}_{2})+\text{f}_{\varepsilon}\widetilde{P}^{\prime}(\bm{x}_{1},\bm{x}_{2})$
@@
Note that $P(t_{\alpha},t_{\beta},t_{\gamma})=p_{\alpha}(t_{\alpha})p_{\beta}(t_{\beta})p_{\gamma}(t_{\gamma})$ is implied by the last constraint. Furthermore, this constraint is coming from the network structure, which implies $P(a_{2}|0,t_{\beta},t_{\gamma})=P(a_{2}|1,t_{\beta},t_{\gamma})$. Moreover, note that here $p_{\alpha}(t),p_{\beta}(t),p_{\gamma}(t)$ correspond to the PTC distribution $\widetilde{P}$ and not the observed distribution $P$, therefore we do not know their values and we consider them as additional variables. Note that the quadratic constraints in $(ii)$ are not convex, i.e. if we write these constraints in the form of $x^{T}Qx$, the matrix $Q$ is not positive semidefinite.
@@
First notice that $p_{\alpha}(t_{\alpha}=0)=\frac{1}{2}\big{(}1+\sqrt{|\frac{\widetilde{E}_{b_{1}}\widetilde{E}_{c_{1}}}{\widetilde{E}_{a_{1}}}|}\big{)}$. Where, $\widetilde{E}_{x_{1}}=\widetilde{P}(x_{1}=0)-\widetilde{P}(x_{1}=1)$. It is straightforward to see that the total variation distance leads to the conclusion that $|\widetilde{E}_{x_{1}}-E_{x_{1}}|\leq 2\text{f}_{\varepsilon}$. In fact, a slightly better bound can be obtained by using the fact that $\widetilde{P}$ is in the PTC slice, see Section D.2 below,
@@ D.1 Collection of LPs (the grid technique for $p_{\xi}(t_{\xi})$)
Finally, note that the LP has to be satisfied for some values $p_{\xi}(t_{\xi})$ of the token probabilities in the intervals $l_{\alpha},u_{\alpha}]$. In practice, it is convenient to divide each interval into smaller intervals $l_{\alpha}+k\frac{u_{\alpha}-l_{\alpha}}{M},l_{\alpha}+(k+1)\frac{u_{\alpha}-l_{\alpha}}{M}]$, and verify that no solution can be found within any of the $M^{3}$ possible combinations of intervals. This gives a better result as compared to the harsher relaxation in Eqs. (113,114). In particular, for the figures 7,8,12,13, and 14 we used $M=16$, while to obtain the best bounds $\omega=0.544\%,\omega^{\prime}=0.504\%$ white noise and $\varepsilon=0.24\%$ total-variation distance we went to $M=128$.
@@ D.2 Development of a tighter bound on $\widetilde{E}_{x_{1}}$
$P(\bm{x}_{1},\bm{\xi}\in\Lambda)=P(\bm{x}_{1};\bm{\xi}\in\Upsilon)+(\text{f}_{\varepsilon}-\varepsilon)P^{\prime\prime}(\bm{x}_{1})$
$\widetilde{P}(\bm{x}_{1})=P(\bm{x}_{1};\bm{\xi}\in\Upsilon)+\text{f}_{\varepsilon}\widetilde{P^{\prime}}(\bm{x}_{1})$
@@ Appendix E Other noise models
[…]
Note that here for the resulting distribution we have $\text{Pr}(a_{1}\oplus b_{1}\oplus c_{1}=1)=\frac{1}{2}\left(1+(1-\omega^{\prime})^{3}\right)$. In Fig.  13(a), we depict the robustness to measurement white noise of the TBSM distributions with $\lambda_{0}^{2}=0.22$. The maximal amount of noise for which we can prove that nonlocality of the distribution is given by $\omega^{\prime}=0.504\%$ for the case $\phi_{u}=0.42$ and $\phi_{w}=0$.
@@
(a) White noise at the measurement
@@
where $E_{\textup{\,\o \,}}$ is the element associated with the no-click event. The probability distribution obtained with such measurements has five outcomes for each party and is not readily analyzed with the results presented in this paper. However, a trick commonly used for device-independent quantum key distribution is to relabel the no-click outcome to one of the four ”good” outcomes restoring output cardinality.
We find that the optimal strategy (for noise robustness) is for each party to relabel the no-click outcome to  ø $\textup{\,\o \,}\to(1,0)$, which corresponds to the POVM
@@
For this strategy, we prove the nonlocality of the resulting distribution for up to a no-click probability of $p=0.7\%$, see Fig.  14(a).
[…]
@@
[…]
@@
[…]
$K_{1}$ $|\displaystyle=\sqrt{1-\eta}\outerproduct{0}{1}$ (122)
@@
Recall that for $\phi_{w}=0$ the original projectors for $x_{1}=1$ read $|\Pi_{00}=\outerproduct{00}{00}$ and $|\Pi_{01}=\outerproduct{11}{11}$, while the lossy POVM elements read
@@ Appendix F Linear Program for the total variation distance
We will prove that all distributions in $\mathcal{B}_{\varepsilon}(\bar{P})$ are nonlocal by contradiction. Let us assume that the ball contains some triangle-local distribution $P$, then from Result 3 there must exist a local PTC distribution $\widetilde{P}$ with $\delta(\widetilde{P}(\bm{x}_{1},\bm{x}_{2}),P(\bm{x}_{1},\bm{x}_{2}))\leq\text{f}_{\varepsilon}$ and $\delta(\widetilde{P}(\bm{x}_{2}),P(\bm{x}_{2}))=0$. By triangle inequality, $\widetilde{P}$ must also satisfy
@@
$\bar{P}(\bm{x}_{1},\bm{x}_{2})=\underbrace{\sum_{\bm{t}}P_{\scriptscriptstyle PTC}(\bm{x}_{1}|\bm{t})P(\bm{x}_{2},\bm{t})}_{\widetilde{P}(\bm{x}_{1}\bm{x}_{2})}-(\varepsilon+\text{f}_{\varepsilon})\bar{P}^{\prime}(\bm{x}_{1},\bm{x}_{2})+(\varepsilon+\text{f}_{\varepsilon})\widetilde{P}^{\prime}(\bm{x}_{1},\bm{x}_{2}).$ (146)
@@
$\bar{P}(\bm{x}_{2})=\widetilde{P}(\bm{x}_{2})-(\varepsilon+\text{f}_{\varepsilon})\bar{P}^{\prime}(\bm{x}_{2})+(\varepsilon+\text{f}_{\varepsilon})\widetilde{P}^{\prime}(\bm{x}_{2}).$ (147)
@@
$\delta(\widetilde{P}^{\prime}(\bm{x}_{2}),\bar{P}^{\prime}(\bm{x}_{2}))=\frac{1}{\varepsilon+\text{f}_{\varepsilon}}\delta(\widetilde{P}(\bm{x}_{2}),\bar{P}(\bm{x}_{2}))\leq\frac{\varepsilon}{\varepsilon+\text{f}_{\varepsilon}}:=\alpha.$ (148)
Note that as $2.5\varepsilon\leq\text{f}_{\varepsilon}\leq 3\varepsilon$, we have $1/4\leq\alpha\leq 2/7$, moreover for small values of $\varepsilon$, we have $\text{f}_{\varepsilon}\approx 2.5\varepsilon$ and $\alpha\approx 2/7$. Again by Lemma 1, we have
$(4.iii^{\prime})\,\sum_{\bm{x}_{1}}\widetilde{P}^{\prime}(\bm{x}_{1},\bm{x}_{2})-\sum_{\bm{x}_{1}}\bar{P}^{\prime}(\bm{x}_{1},\bm{x}_{2})+\alpha\widetilde{P}^{\prime\prime}(\bm{x}_{2})-\alpha\bar{P}^{\prime\prime}(\bm{x}_{2})=0.$ (149)
@@
$(ii)\ p_{\alpha}(0)\sum_{b_{2},c_{2}}{P(a_{2},b_{2},c_{2},0,t_{\beta},t_{\gamma})}-p_{\alpha}(1)\sum_{b_{2},c_{2}}P(a_{2},b_{2},c_{2},1,t_{\beta},t_{\gamma})=0\quad\circlearrowleft.$ (150)
@@
Notice that $p_{\alpha}(t_{\alpha}=0)=\frac{1}{2}\big{(}1+\sqrt{|\frac{\widetilde{E}_{b_{1}}\widetilde{E}_{c_{1}}}{\widetilde{E}_{a_{1}}}|}\big{)}$. Where, $\widetilde{E}_{x_{1}}=\widetilde{P}(x_{1}=0)-\widetilde{P}(x_{1}=1)$. It is straightforward to see that the total variation distance $\delta(\widetilde{P},\bar{P})\leq\text{f}_{\varepsilon}+\varepsilon$, leads to the conclusion that $|\widetilde{E}_{x_{1}}-\bar{E}_{x_{1}}|\leq 2(\text{f}_{\varepsilon}+\varepsilon)$. Where $\bar{E}_{x_{1}}=\bar{P}(x_{1}=0)-\bar{P}(x_{1}=1)$ is computed from the target PTC distribution $\bar{P}$. Hence, if all the correlators for the target distribution satisfy $|\bar{E}_{x_{1}}|\geq 2(\text{f}_{\varepsilon}+\varepsilon)$ one can use Eq. (5) to obtain
$\begin{split}l_{\alpha}:=\frac{1}{2}\left(1+\sqrt{\frac{(|\bar{E}_{b_{1}}|-2(\text{f}_{\varepsilon}+\varepsilon))(|\bar{E}_{c_{1}}|-2(\text{f}_{\varepsilon}+\varepsilon))}{|\bar{E}_{a_{1}}|+2(\text{f}_{\varepsilon}+\varepsilon)}}\right)\leq\ p_{\alpha}(0)\ \\ p_{\alpha}(0)\leq\frac{1}{2}\left(1+\sqrt{\frac{(|\bar{E}_{b_{1}}|+2(\text{f}_{\varepsilon}+\varepsilon))(|\bar{E}_{c_{1}}|+2(\text{f}_{\varepsilon}+\varepsilon))}{|\bar{E}_{a_{1}}|-2(\text{f}_{\varepsilon}+\varepsilon)}}\right):=u_{\alpha}\quad\circlearrowleft.\end{split}$ (151)
@@
$\begin{split}\ l_{\alpha}\sum_{b_{2},c_{2}}{P(a_{2},b_{2},c_{2},0,t_{\beta},t_{\gamma})}-(1-l_{\alpha})\sum_{b_{2},c_{2}}P(a_{2},b_{2},c_{2},1,t_{\beta},t_{\gamma})\leq 0\\ \ u_{\alpha}\sum_{b_{2},c_{2}}{P(a_{2},b_{2},c_{2},0,t_{\beta},t_{\gamma})}-(1-u_{\alpha})\sum_{b_{2},c_{2}}P(a_{2},b_{2},c_{2},1,t_{\beta},t_{\gamma})\geq 0.\end{split}$ (152)
Combining Equations 146, 149, and 152, yields a set of linear constraints that should be satisfied if there exists any local distribution inside the ball $\mathcal{B}_{\varepsilon}(\bar{P})$. Therefore, the infeasibility of this LP proves the non-locality of all of the points in $\mathcal{B}_{\varepsilon}(\bar{P})$. (see Fig.  8). Again, to find tighter bounds, we use the grid technique.